OpenStudy (anonymous):

I would like to create a rectangular orchid garden that abuts my house so that the house itself forms the northern boundary. The fencing for the southern boundary costs $4 per foot, and the fencing for the east and west sides costs $2 per foot. If I have a budget of $80 for the project, what are the dimensions of the garden with the largest area I can enclose? HINT [See Example 2.] … ft(smaller value) x … ft (larger value) … ft2

OpenStudy (anonymous):



OpenStudy (anonymous):

have you made up an equation using the 80, and 2 and 4 dollars yet?

OpenStudy (anonymous):

I did 4x+4y=80

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

yep, good. now we need an equation for area..

OpenStudy (anonymous):

Would it just be the equation of a regular rectangle?

OpenStudy (anonymous):

A = wh ?

OpenStudy (anonymous):

yep, so with this question we want to maximize the area, so that means we will want w and h to be the same value

OpenStudy (anonymous):

so going back to this equation: 4x+4y=80 we want x and y to be the same value, to maximize area with this particular problem, so we can do x=y thus 4x+4x=80 now solve for x

OpenStudy (anonymous):

x=20-y ?

OpenStudy (anonymous):

not quite... we had 4x+4y=80 and to maximize the area of this problem, we want x and y to be equal lengths because they both have a 4 as a coefficient. so we can say y=x so this: 4x+4y=80 becomes 4x+4x=80 now just solve for x

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

so A = 100

OpenStudy (anonymous):

Yay :) It's right!

OpenStudy (anonymous):

Thank you! :) 2 more questions to go :p

OpenStudy (anonymous):

yep ^_^ good job just remember when ever your equation looks like ours did where x and y both had a 4 infront of them, then you have x=y